The following production and average cost data for two levels of monthly production volume have been supplied by a company that produces a single product:

Production volume 1,000 units 3,000 units
Direct materials $ 13.20 per unit $ 13.20 per unit
Direct labor $ 14.50 per unit $ 14.50 per unit
Manufacturing overhead $ 65.40 per unit $ 29.40 per unit


The best estimate of the total variable manufacturing cost per unit is:

Multiple Choice

  • $39.10

  • $27.70

  • $11.40

  • $13.20
Answer #1

Overhead cost for 1000 units = (1000*65.40) = 65,400

Overhead cost for 3000 units = (3000*29.40) = 88,200

Variable Manufacturing Overhead = Change in cost/Change in output

= (88,200-65,400)/(3000-1000)

= 11.40 per unit

Total Variable Manufacturing cost = Direct Materials + Direct Labor + Vmoh

= 13.20+14.50+11.40

= 39.10